A point moves along the path Y= (X/3) with a constant speed of Sm/s. What are the X and Y components of velocity when X=3m? What is the acceleration of point when x=3m!


1
Expert's answer
2021-04-12T01:47:46-0400

"y= \frac{x}{3}"

Differentiating "\frac{dy}{dt}=3 \frac{dx}{dt}"

"\frac{dy}{dt}=v_y; \frac{dx}{dt}=v_x"

Also from "x^2+y^2=s^2" we know that "\sqrt{v_x^2+v_y^2}=s= const"

At x=3, "\frac{dy}{dt}=3 \frac{dx}{dt} \implies v_y=3v_x"

1) "v_x^2+9v_x^2=s"

"10v_x^2=s"

"v_x= \sqrt{\frac{s}{10}}m/s"

"v_y= 3\sqrt{\frac{s}{10}} m/s"

And acceleration, "a= \frac{v^2}{r}"

Radius of curvature ,"r=\frac{(1+ (\frac{dy}{dx})^2)^3}{\frac{d^2y}{dx^2}}"

r="\frac{(1+3^2)^3}{0}=0"

"a= \frac{s^2}{0}=0 m/s^2"